OpenStudy (anonymous):

anyone in connections 9th grade algebra 1b

OpenStudy (haseeb96):

post your question?

OpenStudy (anonymous):

I am not in 9th grade but i wanna give it a try

OpenStudy (anonymous):

ya i can't do that

OpenStudy (anonymous):

I am gonna make estimated guess and say its either B or D

OpenStudy (anonymous):

for First one

OpenStudy (haseeb96):

sqrt t +9=16 sqrt t =16-9 sqrt t= 7 taking square root on both sides t = 7^2 t=49

OpenStudy (haseeb96):

sqrt( 6g-23)=sqrt(12-g) taking square root on boht sides 6g-23=12-g 6g +g =12+23 7g=35 g=5